# An Empirical Analysis on the Use and Reporting of National Security Letters

### v3

@@
Abstract
National Security Letters (NSLs) are a form of legal process that empowers parts of the United States federal government to request certain information for national security purposes. Authorized under five distinct statutory provisions, NSLs are similar to administrative subpoenas and can be issued directly by elements of the executive branch without requiring prior approval from a court or grand jury. Importantly, NSLs authorize the imposition of nondisclosure orders (aka “gag orders”) on the receiving party. Controversy about potential abuses of this authority has driven a range of legal and policy discussions. To address these concerns, both the public sector (through Congressional reporting requirements) and the private sector (through corporate “transparency reports”) have sought to document the usage of NSLs in aggregated form. In addition, some NSLs are now made public due to changes in the law that placed conditions on the nondisclosure requirements of NSLs. However, each of these disclosures is limited in scope, time, and kind. It remains unclear to what extent the heterogeneous data from different sources can be combined to draw meaningful conclusions about the usage of NSLs.
In this paper, we attempt to improve this state of affairs by consolidating the available data around NSLs from various sources and analyzing it. Our data collection allows us to answer two questions: (1) what can the public effectively learn from the reported data and does this information suffice to assess the NSL usage? and (2) how accessible is this data collection? On the one hand, we show that after consolidating and processing data, longitudinal trends in the usage of NSLs can be observed. For instance, we find a significant increase in NSL requests for non-US persons starting from 2015. The observed trends suggest that the current transparency mechanisms are viable safeguards against the excessive use of NSLs. On the other hand, we find that aggregating and normalizing the data is a challenging task, and one of our main contributions is producing a normalized and accessible data set and making it available for future research. The various data sources pose major challenges as they all lack a standardized and machine-readable format. Releasing transparency reports and NSLs, after their gag order was lifted, is optional for companies and only done consistently by a few. While the government data is comparatively straightforward to find, as the publishing entities are known, processing them is challenging. Indeed, the format of the reported data changed repeatedly and the publication as scanned PDFs with continuous text requires manual reviewing, parsing, and validating of the data. We even find inconsistencies within and across data sources. Overall, the laborious data collection process hinders external and internal auditing efforts and demonstrates the need for a unified and more usable dataset for NSLs. We advocate for publishing NSL statistics in a standardized and easily processable format to facilitate third-party data analysis that ensures that Congress and the public can satisfy their role of overseeing the use of NSLs by government agencies.
1 Introduction
National Security Letters, or NSLs, are legal requests issued by US government agencies (predominantly the FBI) to acquire information for investigating national security issues. As statutorily authorized, named agencies can request information about individuals from a broad array of third-party sectors, including financial institutions, communication providers, and credit agencies without requiring approval from the courts or grand jury.
As originally constituted, NSLs were a powerful investigative tool with little judicial oversight, relaxed reporting requirements, and nondisclosure provisions preventing companies from reporting the reception of NSLs or their content. Perhaps due to these properties, NSL usage grew rapidly [26], and lower courts observed that the NSL practices might conflict with the First Amendment and the separation of powers [43, 42]. Subsequently, Congress attempted to increase transparency and congressional oversight for NSLs, starting with the USA Patriot Improvement and Reauthorization Act statutes in 2006. After Edward Snowden’s disclosures concerning government surveillance activities in 2013 [29], companies further pushed back against NSL, fighting for more transparency, and obtained permission to publish their own NSL statistics (albeit quantized in buckets of size 250 or more to obscure precise analysis) [20, 15].
While there long have been efforts to create a more modern and transparent government, these developments have not yet been adopted for NSL statistics. Starting in 2009, the Open Government Initiative aims to “[i]mprove public access to government data, research, and information to enhance transparency, accountability, and equitable outcomes” [30] and Data.gov [25] hosts over 290,000 raw data sets. In comparison, the NSL data does not follow any of the best practices for publishing structured, machine-readable government data discussed by Joe Calandrino and Harlan Yu in a series of blog posts [23]. The NSL information is published by different government bodies, in PDFs that require manual labor to parse, and with frequently changing data formats. This lack of a universal data format creates an obstacle to analysis, as it requires a significant amount of labor to acquire, extract, and normalize the information scattered across different reports. Perhaps, security concerns disincentivize agencies from releasing too fine-grained NSL statistics. However, while the current unstructured data publishing does not preclude analysis, the intricate and cumbersome process of doing so hinders the public from fulfilling their role of overseeing surveillance authorities. Despite that there has been significant discussion in the legal community about the constitutionality of NSL provisions in the context of First and Fourth Amendment issues [41, 28, 24, 32, 22, 27], thus far, we are unaware of any comprehensive analysis of NSL reports. Similar to government data, information published by companies shows a similar lack of structure, especially across organizations.
In this paper, we create a consolidated dataset from disparate NSL data sources. Using a combination of manual cleaning and automated scripts, we extract NSL-related information from a broad array of reports from both the public and private sectors. This data enables us to perform an empirical analysis on the use of NSLs over time, as well as to identify both corroboration and inconsistencies across data sources. We publish the resulting data set to allow other entities to assess the use of NSLs.
Overall, our paper makes three main contributions:
@@ 1 Introduction
  • First, we present the first dataset that consolidates heterogeneous data from government statistics, transparency reports, and published NSLs. With a combination of manual efforts and automated scripts, we produce a normalized dataset that enables comparisons across data sources. We make our scripts and data available to the public to enable future research.
@@
  • Second, we observe that the analysis of this data enables insights about the use and reporting of NSLs over time, including the following: NSL requests for non-US persons grew significantly between 2010 and 2015, passing the number of requests for US persons, and remaining popular until today, with significant spikes in frequency in 2015 and 2019. Furthermore, transparency reports show that telecommunication companies (e.g., AT&T, T-Mobile, and Verizon) receive the largest number of NSLs among the reporting companies. Our cross-comparison of transparency reports with the number of NSL requests reported by the government appears to be consistent. Finally, the metadata of published NSLs (after their gag order was lifted) lets us infer their time under disclosure—a metric not reported by the government—which appears to be decreasing since 2010, in response to legal reforms. These analyses suggest that the current reporting mechanisms are effective in preventing the abuse of NSLs.
  • Third, we document the challenging task of processing data from various sources due to the absence of standardized and machine-readable formats. This laborious process significantly hinders external and internal auditing efforts. Indeed, in our analysis, we have found various inconsistencies both within and across data sources. Our work demonstrates the need for a unified and more usable dataset for NSLs, which would ultimately lead to better utilization and appreciation of the NSL data available as well as a government with increased transparency.
2 NSL: A Brief Background
Currently, five statutory provisions authorize government agencies to issue NSLs: the Right to Financial Privacy Act (RFPA) [2]; the Electronic Communications Privacy Act (ECPA) [4]; the National Security Act (NSA) [1]; the Fair Credit Reporting Act [7]; and the USA PATRIOT Act [9] (the last of which being the first to codify the term “National Security Letter” explicitly).
@@ 2 NSL: A Brief Background
In this section, we start by giving a short overview of these five NSL statutes, followed by a summary of efforts to improve transparency around the use of NSLs (notably two amendments to the five NSL statutes that introduced reporting mandates).
2.1 The Five NSL Statutes
@@
The Right to Financial Privacy Act (RFPA) [2] was the first statute used to introduce NSL authorities. As part of a 1986 amendment, the FBI was granted the right to request access to business records from financial institutions (amended in 2003 to include a broader range of organizations). During the same period of time, the second NSL-granting statute, the Electronic Communications Privacy Act (ECPA) [4] was enacted. It provided access (via 18 U.S.C § 2709) to business records (i.e., name, address, length of service, and toll records) of wire or electronic communication service providers for counterintelligence purposes.
@@
In the 1990s, two more statutes were enacted granting further NSL authorities. In 1994, the National Security Act (NSA) was amended to include a procedure for any authorized agency to request a broad array of business records from various organizations for investigating potential document leaks from government employees (codified at 50 U.S.C § 3162 [6]). Later in 1996, the Fair Credit Reporting Act (FCRA) was amended to incorporate the fourth statutory provision for NSLs, which authorized the FBI to access credit agency records in service of national security investigations.
@@
Finally, the PATRIOT Act [9] is the fifth and last NSL-granting statute, which was enacted in 2001 as a response to the terrorist attacks of September 11 [21]. Through amending FCRA, the PATRIOT Act introduced a new procedure for government agencies to access consumer reports from credit report agencies for counterintelligence and counterterrorism investigations. In addition, it made substantial amendments to three of the four existing NSL statutes (RFPA, ECPA, and FCRA), expanding the scope of NSLs and simplifying the administrative approval requirements.
In addition to providing access to sensitive records from an array of private sector institutions, the five NSL statutes initially had no mandatory judicial review, indefinite nondisclosure mandates, and few reporting requirements, making NSLs a powerful tool with limited supervision. Due to concerns from courts and the public, various efforts have been made to increase the transparency of NSL usage and congressional oversight by introducing reporting mandates and weakening nondisclosure requirements, which we discuss below.
2.2 Subsequent Amendments
2.2.1 RFPA amendment in 2003
2.2.2 PATRIOT Act amendments
In 2006, the 109th Congress amended the USA PATRIOT Act with two statues, in part as a response to earlier judicial reactions to the USA PATRIOT Act [32]. Specifically, two court cases Doe v. Gonzales [42] and Doe v. Ashcroft [43]111 The plaintiff’s name “John Doe” is used in both cases because the gag order prohibited the disclosure of the plaintiff’s identity. raised First and Fourth Amendment issues with NSLs connected to the non-disclosure provisions and the absence of judicial oversight [26]. In the case of Doe v. Ashcroft, the district court held that NSLs violated the Fourth Amendment because they authorized “coercive searches effectively immune from any judicial process” [45]. Moreover, the court held that the nondisclosure provisions unconstitutionally restrict free speech as they can prohibit disclosure without providing judicial means to challenge a ban or achieve eventual relief [45, 44].
As a reaction to both cases, and while Doe v. Ashcroft was on appeal at the Second Circuit, the 109th Congress amended the USA PATRIOT Act with the USA PATRIOT Improvement and Reauthorization Act [13] and the USA PATRIOT Act Reauthorization Amendments Act [12]. Collectively, besides adding penalties for non-compliance with NSLs or their nondisclosure requirements [11, 10], these two statutes introduced judicial review on the use of NSLs and weakened the nondisclosure requirements. Specifically, the amendments require that the Foreign Intelligence Surveillance Court (FISC) approves NSL requests and adds control mechanisms and a process to ease nondisclosure provisions [11, 10]. Regarding Doe v. Ashcroft, the Second Circuit ruled in 2008 that the FBI needs to certify that disclosure of an NSL would lead to statutorily enumerated harms to justify non-disclosure provisions [31]. Last, the amendments call for an audit on use of the NSLs from the Office of the Inspector General (referred to as the OGI reports).
2.2.3 USA FREEDOM Act
In 2015, Congress enacted the USA FREEDOM Act [19], which revised the nondisclosure requirement, introduced an additional reporting requirement, and expressly limited the use of NSL to specifically identified information. Notably, the additional reporting requirement and restrictions on the use of NSLs were direct responses to the bulk metadata collection practice of the NSA [39].
2.3 Reporting and Nondisclosure Mandates of NSLs
@@
Initially, the five NSL statutes had no mandatory judicial review, relaxed reporting requirements, and strict nondisclosure mandates [^2] As an example, the original ECPA and FCRA both prohibited the disclosure to any person. . Given the broad range of data accessible via NSLs, this has unsurprisingly sparked concerns both in the public and, in response to challenges, in the courts [26]. Over time, partially fueled by unfavorable judicial reactions (e.g., Doe v. Gonzales [42]) and various public events (e.g., the Snowden disclosures in 2013 [29]), several efforts have been made to increase the transparency of NSL usage [41, 26]. We highlight the two legislative efforts that are most relevant to this paper.
@@
In 2006, Congress amended the NSL statutes with the USA PATRIOT Improvement and Reauthorization Act [13] and the USA PATRIOT Act Reauthorization Amendments Act [12], establishing a process for judicial review, weakening nondisclosure requirements, and introducing additional reporting requirements. Notably, the USA PATRIOT Improvement and Reauthorization Act added two unclassified reporting requirements. Section 119 of the Act initiated an audit by the Office of Inspector General of the Department of Justice to report NSL usage to Congress (“the OIG reports”). However, these reports were each singular occurrences, and the OIG is not expected to publish any further reports unless there is another mandate from Congress. Second, Section 118 required the Attorney General of the Department of Justice to submit annual unclassified statistics about NSLs targeting US persons to Congress [14]. These statistics are published as part of the Foreign Intelligence Surveillance Act reports, hence referred to as FISA reports.
@@
In 2015, the USA FREEDOM Act [19] mandated additional reporting, and expressly limited the use of NSLs to explicitly specified information. Notably, it mandates that the Office of the Director of National Intelligence (ODNI) should publish a report that details both the number of NSLs issued, and the number of requests for information contained in those NSLs for the past year (in addition to the FISA reports) [19]. This is published as part of the National Intelligence’s Annual Statistical Transparency Report (ASTR).
@@
Finally, in addition to adding reporting mandates and limiting the scope of NSLs, the USA FREEDOM Act introduced three major changes that relaxed the nondisclosure requirements. First, it allowed companies to report the total number of NSLs received (and customers covered by those NSLs) quantized into bands. [^3] Public records [26, 17] suggested that this part of the USA Freedom Act was inspired by the pre-existing voluntary agreement reached in 2014 between several technology companies and the DOJ [15]. In practice, a range of companies (particularly in the technology and communications sectors) take advantage of this permission and publish a rough estimate of the number of NSLs they receive as part of their annual transparency reports. We refer to such reports collectively as “transparency reports”. Moreover, the USA FREEDOM Act codified a procedure (18 U.S.C. § 3511) that allows companies to request judicial review of the nondisclosure orders (commonly known as “reciprocal notice”). Lastly, it required the Attorney General to adopt procedures to review nondisclosure orders at appropriate intervals. If disclosure is no longer believed to lead to harms listed in 18 U.S.C. § 2709(c)[5], then the FBI should terminate the nondisclosure order and notify the company. Some companies publish the content of such NSLs (i.e., after the termination of nondisclosure requirements) which can provide useful metadata for analyzing past NSL issuance. We refer to such publications collectively as “company NSLs”.
@@
  • OIG, FISA and ASTR reports (collectively referred to as “Government Reports” in Section 3.1) mandated by the USA PATRIOT Improvement and Reauthorization Act and the USA FREEDOM Act
@@
Understanding what the public can learn from these publicly available NSL reports, and the limitations of each data source is the central focus of our work.
3 Data Collection and Trends

### v4

@@
Abstract.
Government investigatory and surveillance powers are important tools for examining crime and protecting public safety. However, since these tools must be employed in secret, it can be challenging to identify abuses or changes in use that could be of significant public interest. In this paper, we evaluate this phenomenon in the context of National Security Letters (NSLs). NSLs are a form of legal process that empowers parts of the United States federal government to request certain pieces of information for national security purposes. After initial concerns about the lack of public oversight, Congress worked to increase transparency by mandating government agencies to publish aggregated statistics on the NSL usage and by allowing the private sector to report information on NSLs in transparency reports. The implicit goal is that these transparency mechanisms should deter large-scale abuse by making it visible. We evaluate how well these mechanisms work by carefully analyzing the full range of publicly available data related to NSL use. Our findings suggest that they may not lead to the desired public scrutiny as we find published information requires significant manual effort to collect and parse data due to the lack of structure and context. Moreover, we discovered mistakes (subsequently fixed after our reporting to the ODNI), which suggests a lack of active auditing. Taken together, our case study of NSLs provides insights and suggestions for the successful construction of transparency mechanisms that enable effective public auditing.
empirical analysis, national security letters, NSL, transparency, auditability
ccs: Applied computing Lawccs: Applied computing E-governmentjournalyear: 2025copyright: ccconference: Symposium on Computer Science and Law; March 25–27, 2025; München, Germanybooktitle: Symposium on Computer Science and Law (CSLAW ’25), March 25–27, 2025, München, Germanydoi: 10.1145/3709025.3712209isbn: 979-8-4007-1421-4/25/03
1. Introduction
Intelligence and surveillance statutes such as the Foreign Intelligence Surveillance Act of 1978 and the USA PATRIOT Act of 2001 provision powerful legal tools that enable the government to collect a broad array of information for investigating crimes and protecting public safety. Due to the sensitive nature of such investigations, information about the content, target, or even existence of such requests may be security critical. However, without any public information on their use, there is no oversight to ensure proper use and deter abuse. Indeed, incidents of abuse have been documented in the past, such as the NSA’s bulk collection of customers’ telecommunication records (guardian-snowden, ). One approach to address these concerns is increased transparency—either via statutory requirements for disclosure or documentation (e.g., mandated disclosure of wiretaps or aggregated data about the use of NSLs) or via voluntary reporting by the private sector (e.g., in transparency reports). The implicit reasoning is that even without revealing the specifics of individual operations, large-scale abuses will become evident under careful public scrutiny, and will thus deter such behavior.
However, the effectiveness of this approach hinges upon two assumptions: first, that such disclosures contain sufficient data, documentation and context that an unprivileged third-party could easily audit them for compliance, and second, that there are third-parties with the funding, focus and expertise necessary to do such work on a regular basis. In this work, we investigate the extent to which these two assumptions hold in the context of National Security Letters (NSLs).
We show that both assumptions may be too strong in practice. Although both the public and private sector have sought to provide transparency by documenting NSL usage in aggregate, we document the need for better curation and documentation in this data, as evidenced by the significant amount of manual effort required to collect, parse, and process it. Additionally, we discovered data discrepancies that were subsequently fixed after reporting to ODNI, suggesting the lack of active auditing and checking from the public. We discussed with the former and current Chief, ODNI Civil Liberties, Privacy, and Transparency Office to validate our findings and understand current operational challenges as well as potential future improvements.
Altogether, our work evaluates existing transparency mechanisms for NSLs. We provide empirical data by combining public resources and insights into operational challenges and real-world constraints that future work can consider when exploring new systems and approaches for accountability. We suggest improvements to address some challenges without changing the balance between national security and transparency that the Congress decided on for NSLs.
Concretely, our paper makes the following contributions:
@@ 1. Introduction
  • We present the first dataset that consolidates heterogeneous NSL data from government statistics, transparency reports, and published NSLs. With a combination of manual effort and automated scripts, we produce a normalized dataset that enables comparisons across data sources. We make this data available, both to enable future research and allow for easier auditing by the public.
  • As a result of our data collection, we identify and report data inconsistencies present in the datasets to the relevant government agencies, which have since been corrected (fisa-correction-2020, ).
@@
  • We analyze this data and provide insights about the use and reporting of NSLs over time, including the following: NSL requests for non-US persons grew significantly between 2010 and 2015, passing the number of requests for US persons, and remaining popular until today, with significant spikes in frequency in 2015 and 2019. Furthermore, our analysis of public transparency reports suggest that telecommunication companies (e.g., AT&T, T-Mobile, and Verizon) receive the largest number of NSLs among the reporting companies. Our cross-comparison of transparency reports with the number of NSL requests reported by the government appears to be consistent.
  • We provide suggestions for improving the overall environment of NSL reporting in both public and private sectors. We outline changes that can be made to the government data publishing process that can improve the accuracy and security of data without adding undue burden. In addition, we offer suggestions for improved transparency in the private sector for NSL-related data publishing.
2. Related Work
There are four main lines of research that are related to our work. First, a body of work argues for the importance of transparency in discouraging potential abuse of power by the government, such as judicial power (smith2009kudzu, ) and surveillance activities (smith2012gagged, ; bloch2018exposing, ). These discussions inspired us to empirically evaluate the transparency mechanisms of NSLs.
A second area of research suggests novel mechanisms (frankle2018practical, ; segal2014catching, ; krollsecure, ) that enable transparency and accountability while preserving secrecy. These mechanisms utilize cryptographic primitives and often allow auditing at a fine-grained (e.g., per-case) level without disclosing confidential information or violating privacy. Instead of exploring new mechanisms with optimal transparency and accountability properties, our work evaluates existing mechanisms, operational practices, and challenges. We hope that future work and discussions in Congress can combine our empirical understanding of deployed mechanisms with new ideas—such as (frankle2018practical, ; segal2014catching, ; krollsecure, )—to codify new mechanisms that are practical and effective.
A third research direction related to our work encompasses the empirical analysis of publicly available data. This includes Romanosky et al.’s (romanosky2014empirical, ) analysis of data breach litigations, Beller’s (beller2022401, ) analysis of publicly available notices issued to targets being spied on under the authorization of FISA, and Kesari’s (kesari2023data, ) assessment of identity theft report rates. Our work also takes an empirical approach and utilizes data to gain insights into the state of transparency of NSLs.
Finally, there are various discussions of NSLs in the legal research community. However, most of them are centered around the constitutionality of NSL provisions in the context of First and Fourth Amendment issues (bloch2016process, ; dallal2018speak, ; garlinger2009privacy, ; gorham2008national, ; nieland2006national, ; schrop2017your, ). One exception is the analysis by EPIC (EPICNati14:online, ), which also collects and examines public data. However, their analysis and data collection is limited in depth and scope, and focuses mostly on events between the PATRIOT Act in 2005 and the USA FREEDOM Act in 2015.
3. NSL: A Brief Background
Currently, five statutory provisions authorize government agencies to issue NSLs: the Right to Financial Privacy Act (law:USC-12-Chapter-35-RFPA, ); the Electronic Communications Privacy Act (law:USC-18-2701-ECPA-access-stored-communications, ); the National Security Act (law:PL-80-253-NSA, ); the Fair Credit Reporting Act (law:USC-15-1681-FCRA, ); and the USA PATRIOT Act (law:PL-107-56-PATRIOT, ) (the last of which being the first to codify the term “National Security Letter” explicitly). Together, these statutes allow government agencies—predominantly the FBI—to use NSLs to request metadata for a person (often referred to as “target”) from various companies including electronic communication service providers and credit agencies. A single NSL can contain multiple Requests For Information (ROI)—such as an account identifier or email address—as long as they are in the context of a single investigation. Additionally, the FBI may serve multiple NSLs for the same person under different statutes and to different entities (e.g., to collect information on phone calls, emails, and financial records). The reported usage statistics distinguish NSLs for US persons, non-US persons, and subscriber information. The latter includes NSL letters for both US and non-US persons where the target’s nationality was not known at the time when the NSL letter was issued.
@@ 3. NSL: A Brief Background
In this section, we start by giving a short overview of the five NSL statutes, followed by a summary of efforts to improve transparency around the use of NSLs (notably two amendments to the five NSL statutes that introduced reporting mandates). We refer the interested reader to Appendix A for a more extensive historic background on NSLs.
3.1. The Five NSL Statutes
@@
The Right to Financial Privacy Act (RFPA) (law:USC-12-Chapter-35-RFPA, ) was the first statute used to introduce NSL authorities. As part of a 1986 amendment, the FBI was granted the right to request access to business records from financial institutions (amended in 2003 to include a broader range of organizations). During the same period of time, the second NSL-granting statute, the Electronic Communications Privacy Act (ECPA) (law:USC-18-2701-ECPA-access-stored-communications, ) was enacted. It provided access (via 18 U.S.C § 2709) to business records (i.e., name, address, length of service, and toll records) of wire or electronic communication service providers for counterintelligence purposes.
@@
In the 1990s, two more statutes were enacted granting further NSL authorities. In 1994, the National Security Act (NSA) was amended to include a procedure for any authorized agency to request a broad array of business records from various organizations for investigating potential document leaks from government employees (codified at 50 U.S.C § 3162 (law:USC-50-3162-codified-IAA, )). Later, in 1996, the Fair Credit Reporting Act (FCRA) was amended to incorporate the fourth statutory provision for NSLs, which authorized the FBI to access credit agency records in service of national security investigations.
@@
Finally, the PATRIOT Act (law:PL-107-56-PATRIOT, ) is the fifth and last NSL-granting statute, which was enacted in 2001 as a response to the terrorist attacks of September 11 (wikipedia:patriot-sept-11, ). Through amending FCRA, the PATRIOT Act introduced a new procedure for government agencies to access consumer reports from credit report agencies for counterintelligence and counterterrorism investigations. In addition, it made substantial amendments to three of the four existing NSL statutes (RFPA, ECPA, and FCRA), expanding the scope of NSLs and simplifying the administrative approval requirements.
3.2. Reporting and Nondisclosure Mandates of NSLs
@@
Initially, the five NSL statutes had no clear mechanism for judicial review, relaxed reporting requirements, and strict nondisclosure mandates [^1] As an example, the original ECPA and FCRA both prohibited the disclosure of an NSL request to any person. . Given the broad range of data accessible via NSLs, this unsurprisingly sparked concerns both in the public and, in response to challenges, in the courts (CRS:legalBackground, ). Over time, partially fueled by unfavorable judicial reactions (e.g., Doe v. Gonzales (case:DoeVGonzales, )) and various public events (e.g., the Snowden disclosures in 2013 (guardian-snowden, )), several efforts have been made to increase the transparency of NSL usage (CRS:legalBackground, ). We highlight the two legislative efforts that are most relevant to this paper.
@@
In 2006, Congress amended the NSL statutes with the USA PATRIOT Improvement and Reauthorization Act (law:PL-109-177-PATRIOT-2005, ) and the USA PATRIOT Act Reauthorization Amendments Act (law:PL-109-178-PATRIOT-2005, ). Notably, the USA PATRIOT Improvement and Reauthorization Act added two unclassified reporting requirements. Section 119 of the Act initiated an audit by the Office of Inspector General of the Department of Justice (DOJ) to report NSL usage to Congress (“the OIG reports”). However, these reports were each one-off occurrences, and the OIG is not expected to publish any further reports unless there is another mandate from Congress. Second, Section 118 requires the Attorney General of the DOJ to submit annual unclassified statistics about NSLs targeting US persons to Congress (law:PL-109-177-PATRIOT-2005-section-128, ). These statistics are published as part of the Foreign Intelligence Surveillance Act reports, hence referred to as “FISA reports”.
@@
In 2015, the USA FREEDOM Act (law:PL-114-23-FREEDOM, ) mandated additional reporting, and expressly limited the use of NSLs for explicitly specified information. Notably, it mandates that the Office of the Director of National Intelligence (ODNI) publishes a report that details both the number of NSLs issued, and the number of requests for information contained in those NSLs for the past year (in addition to the FISA reports) (law:PL-114-23-FREEDOM, ). This is published as part of the National Intelligence’s Annual Statistical Transparency Report (ASTR).
@@
Finally, in addition to adding reporting mandates and limiting the scope of NSLs, the USA FREEDOM Act introduced three major changes that relaxed the nondisclosure requirements. First, it allowed companies to report the total number of NSLs received (and customers covered by those NSLs) quantized into bands of 250, 500 or 1000.222Public records (CRS:legalBackground, ; law:HRep-114-109, ) suggested that this part of the USA Freedom Act was inspired by the pre-existing voluntary agreement reached in 2014 between several technology companies and the DOJ (dojagreement:2014, ). In practice, a range of companies (particularly in the technology and communications sectors) take advantage of this permission and publish a rough estimate of the number of NSLs they receive as part of their annual transparency reports. We refer to such reports collectively as “transparency reports”. Moreover, the USA FREEDOM Act codified a procedure (18 U.S.C. § 3511) that allows companies to request judicial review of the nondisclosure orders (commonly known as “reciprocal notice”). Lastly, it required the Attorney General to adopt procedures to review nondisclosure orders at appropriate intervals. If disclosure is no longer believed to lead to harms listed in 18 U.S.C. § 2709(c)(law:USC-18-2709-ECPA-access-to-records, ), then the FBI should terminate the nondisclosure order and notify the company. Some companies have since chosen to publish the content of such NSLs (i.e., after the termination of nondisclosure requirements) which can provide useful metadata for analyzing past NSL issuance. We refer to such publications collectively as “company NSLs”.
@@
  • OIG, FISA and ASTR reports (collectively referred to as “Government Reports” in Section 4.1) mandated by the USA PATRIOT Improvement and Reauthorization Act and USA FREEDOM Act
@@
Understanding what the public can learn from these publicly available NSL reports, how useful the information is to the public, and the limitations of each data source is the central focus of our work.
4. Data Collection and Trends
Figure 1. Overview of three types of data sources: government statistics, transparency reports, and published NSLs
[…]
[Overview of three types of data sources: government statistics, transparency reports, and published NSLs.] Overview of three types of data sources: government statistics, transparency reports, and published NSLs. Type 1 contains ASTR, FISA, and OIG reports from the government. ASTR reports publish the number of NSLs issues as well as the number of total ROIs since 2013. FISA reports include the number of ROIs and targets for US persons since 2005 and additionally publish information on non-US persons and subscriber only information requests starting from 2015. OIG reports published the total number of ROIs, as well as number of US and non-US ROIs between 2003 and 2009, and only the total number of ROIs between 2009 and 2011. Type 2 are transparency reports where companies report the number of NSLs they received in bands of 250, 500, or 1000. Type 3 are published NSLs containing metadata such as file numbers, as well as issuance and publication dates.